OpenStudy (anonymous):

The mass of a sample of sodium bicarbonate is 2.1 kilograms (kg). There are 1,000 grams (g) in 1 kg, and 1 mc020-1.jpg 109 nanograms (ng) in 1 g. What is the mass of this sample in ng?

OpenStudy (anonymous):

Use the factor-label method, if you are familiar with that. When you do the dimensional analysis, if units cancel out, you know you are on the right track.\[1~\text{kg} = 1\;000~\text{g}, \quad 1~\text{g} = 1\times 10^9~\text{ng}\]\[2.1~\text{kg} \times \frac{1\;000~\text{g}}{1~\text{kg}} \times \frac{10^9~\text{ng}}{1~\text{g}}\]\[2.1~\text{kg} = 2.1 \times 10^{12}~\text{ng}\]